# Duality for Fitting’s Multi-valued Modal logic via bitopology and bi-Vietoris coalgebra

### v2

@@ Abstract
Fitting’s Heyting-valued logic and Heyting-valued modal logic have already been studied from an algebraic viewpoint. In addition to algebraic axiomatizations with the completeness of Fitting’s Heyting-valued logic and Heyting-valued modal logic, both topological and coalgebraic dualities have also been developed for algebras of Fitting’s Heyting-valued modal logic. Bitopological methods have recently been employed to investigate duality for Fitting’s Heyting-valued logic. However, the concepts of bitopology and biVietoris coalgebras are conspicuously absent from the development of dualities for Fitting’s many-valued modal logic. With this study, we try to bridge that gap. We develop a bitopological duality for algebras of Fitting’s Heyting-valued modal logic. We construct a bi-Vietoris functor on the category $PBS_{\mathcal{L}}$ of $\mathcal{L}$-valued ($\mathcal{L}$ is a Heyting algebra) pairwise Boolean spaces. Finally, we obtain a dual equivalence between categories of biVietoris coalgebras and algebras of Fitting’s Heyting-valued modal logic. As a result, we conclude that Fitting’s many-valued modal logic is sound and complete with respect to the coalgebras of a biVietoris functor. We discuss the application of this coalgebraic approach to bitopological duality.
@@ 1 Introduction
Algebraic axiomatization of a modified version of Fitting’s Heyting-valued modal logic has already been addressed in [11]. Maruyama [12] proposed Jónsson-Tarski topological duality (see [17, 18, 19]) for $\mathcal{L}$$-\mathcal{ML}$-algebras (algebras of Fitting’s Heyting-valued modal logic). Jónsson-Tarski duality for $\mathcal{L}$$-\mathcal{ML}$-algebras is essentially a $\mathcal{L}$-valued version of Jónsson-Tarski duality for modal algebras.
We aim to construct a bitopological duality for algebras of Fitting’s Heyting- valued modal logic by setting up a notion of $PRBS_{\mathcal{L}}$ as a category of $\mathcal{L}$-valued pairwise Boolean spaces with a binary relation. As a result, natural duality theory for modal algebras is extended in the context of bitopological languages. The main results are bitopological and coalgebraic dualities for $\mathcal{L}$$-\mathcal{ML}$-algebras, where $\mathcal{L}$ is a semi-primal algebra having a bounded lattice reduct. Our general theory extends both the Jónsson-Tarski duality and the Abramsky-Kupke-Kurz-Venema coalgebraic duality [1, 20] in the setting of bitopological language. Furthermore, it introduces a novel coalgebraic duality for $\mathcal{L}$$-\mathcal{ML}$-algebras.
@@
In this study, we seek to demonstrate how coalgebras for the biVietoris functor on the category $PBS_{\mathcal{L}}$ of $\mathcal{L}$-valued version of pairwise Boolean spaces and pairwise continuous maps can be used to understand the semantics of Fitting’s many-valued modal logic.
Our study develops both bitopological and coalgebraic dualities for algebras of Fitting’s many-valued modal logic, building upon previous research [14, 13, 7, 8]. Maruyama et al. [13] developed both topological and coalgebraic dualities for general algebraic structures encompassing algebras of Fitting’s many-valued modal logic. Our research directly contributes to the results of the publication [13]; specifically, we present bitopological versions of those dualities for Fitting’s many-valued modal logic.
@@
In Section 2, we review the fundamental notions of bitopological spaces and algebras of Fitting’s Heyting-valued logic, denoted as $\mathcal{L}$-$\mathcal{VL}$-algebras, and we also discuss the bitopological duality theorem for $\mathcal{L}$-$\mathcal{VL}$-algebras. Section 3 contains our first result, i.e., bitopological duality for Fitting’s Heyting-valued modal logic. In Section 4, we introduce the concept of pairwise Vietoris spaces and construct an endofunctor $V_{\mathcal{L}}^{bi}$ on the category $PBS_{\mathcal{L}}$. Finally, we show the coalgebraic duality for Fitting’s Heyting-valued modal logic. In Section 5, we conclude the paper by outlining several potential future research directions.
@@ 2.2 Fitting’s Heyting-valued logics
Fitting [10] proposed $\mathcal{L}$-valued logics and $\mathcal{L}$-valued modal logics for a finite distributive lattice $\mathcal{L}$(i.e., $\mathcal{L}$ is a Heyting algebra) in 1991. Maruyama [11] introduced algebraic axiomatization of Fitting’s logics. In [11] the author studied Fitting’s Heyting-valued logic and Heyting-valued modal logic without regard for fuzzy truth constants other than 00 and $1$, and added a new operation $T_{\ell}(-)$, $\ell\in\mathcal{L}$. From a logical perspective, $T_{\ell}(p)$ infers that the truth value of a proposition $p$ is $\ell$. The operations of $\mathcal{L}$-valued logic, denoted by $\mathcal{L}$-$\mathcal{VL}$, are $\vee,\wedge,\rightarrow,0,1$ and $T_{\ell}(-)$, $\ell\in\mathcal{L}$, where ,$\vee$,$\wedge$,$\rightarrow$ , ∧ , → are binary operations, 00 and $1$ are nullary operations and $T_{\ell}$ is a unary operation. For $\ell_{1},\ell_{2}\in\mathcal{L}$, $\ell_{1}\rightarrow\ell_{2}$ means the pseudo-complement of $\ell_{1}$ relative to $\ell_{2}$.
@@
$T_{\ell}(\ell^{\prime})=\left\{\begin{array}[]{ll}1&(\ell^{\prime}=\ell)\\ 0&(\ell^{\prime}\neq\ell)\\ \end{array}\right.$
@@
$U_{\ell}(\ell^{\prime})=\left\{\begin{array}[]{ll}1&(\ell^{\prime}\geq\ell)\\ 0&(\ell^{\prime}\ngeq\ell)\\ \end{array}\right.$
@@
The class of all $\mathcal{L}$-$\mathcal{VL}$-algebras forms a variety (in the sense of universal algebra). If =$\mathcal{L}=\{0,1\}$, then $\mathcal{L}$-$\mathcal{VL}$-algebras become Boolean algebras.
@@ 2.3 Bitopological duality for Fitting’s Heyting-valued logic
    For an object $(\mathcal{B},\alpha_{\mathcal{B}})$ in $PBS_{\mathcal{L}}$, define $\mathfrak{F}(\mathcal{B},\alpha_{\mathcal{B}})=(HOM_{PBS_{\mathcal{L}}}((\mathcal{B},\alpha_{\mathcal{B}}),(\mathcal{L},\alpha_{\mathcal{L}})),\vee,\wedge,\rightarrow,T_{\ell}(\ell\in\mathcal{L}),0,1)$, where $\vee,\wedge,\rightarrow,T_{\ell}(\ell\in\mathcal{L}),0,1$ are pointwise operations on the set $HOM_{PBS_{\mathcal{L}}}((\mathcal{B},\alpha_{\mathcal{B}}),(\mathcal{L},\alpha_{\mathcal{L}}))$. The operations 00 and $1$ are regarded as constant functions, with 00 and $1$ being their respective values.
@@ 3 Bitopological duality for Fitting’s many-valued modal logic
$[\mathcal{R}]C=\{p\in P:\mathcal{R}[p]\subseteq C\}$ and $\langle\mathcal{R}\rangle C=\{p\in P:\mathcal{R}[p]\cap C\neq\emptyset\}$.
@@ 3.1 Category
      $\forall\mathcal{C}\in\beta_{1}$, $[\mathcal{R}]\mathcal{C},\langle\mathcal{R}\rangle\mathcal{C}\in\beta_{1}$;
@@
We see that $[\mathcal{R}]Q^{c}=(\langle\mathcal{R}\rangle Q)^{c}$, and $\langle\mathcal{R}\rangle Q^{c}=([\mathcal{R}]Q)^{c}$. Since $\beta_{2}=\{Q^{c}:Q\in\beta_{1}\}$, hence if the relation $\mathcal{R}$ satisfies condition $(ii)$ that is given in the object part of Definition 10, then $[\mathcal{R}]Q,\langle\mathcal{R}\rangle Q\in\beta_{2}$, $\forall Q\in\beta_{2}$.
@@ 3.2 Functors
Definition 9 shows that $(HOM_{\mathcal{VA}_{\mathcal{L}}}(\mathcal{A},\mathcal{L}),\tau_{1},\tau_{2},\alpha_{\mathcal{A}})$ is an object in $PBS_{\mathcal{L}}$. Thus, it suffices to demonstrate $\mathcal{R}_{\Box}$ satisfies the requirements given in the object section of Definition 10. First, we ensure that for any $\mathcal{W}\in HOM_{\mathcal{VA}_{\mathcal{L}}}(\mathcal{A},\mathcal{L})$, $\mathcal{R}_{\Box}[\mathcal{W}]\in\delta_{1}\cup\delta_{2}$. Let $\mathcal{U}\not\in\mathcal{R}_{\Box}[\mathcal{W}]$. Then by Definition 6, there is an element $a\in\mathcal{A}$ such that there is $L_{1}\in\mathcal{L}$, for which $\mathcal{W}(\Box a)\geq L_{1}$ but $\mathcal{U}(a)\not\geq L_{1}$. It follows that $\mathcal{U}\in\langle\neg U_{L_{1}}(a)\rangle\in\tau_{2}$ and $\mathcal{R}_{\Box}[\mathcal{W}]\cap\langle\neg U_{L_{1}}(a)\rangle=\emptyset$ i.e., $\langle\neg U_{L_{1}}(a)\rangle\subseteq(\mathcal{R}_{\Box}[\mathcal{W}])^{c}$. Hence, $\mathcal{U}\not\in\overline{\mathcal{R}_{\Box}[\mathcal{W}]}^{\tau_{2}}$, where $\overline{\mathcal{R}_{\Box}[\mathcal{W}]}^{\tau_{2}}$ denotes the closure of $\mathcal{R}_{\Box}[\mathcal{W}]$ in $(HOM_{\mathcal{VA}_{\mathcal{L}}}(\mathcal{A},\mathcal{L}),\tau_{2})$ Equivalently, we have $\overline{\mathcal{R}_{\Box}[\mathcal{W}]}^{\tau_{2}}\subset\mathcal{R}_{\Box}[\mathcal{W}]$. Therefore, $\mathcal{R}_{\Box}[\mathcal{W}]$ is $\tau_{2}$-closed. Since $(HOM_{\mathcal{VA}_{\mathcal{L}}}(\mathcal{A},\mathcal{L}),\tau_{1},\tau_{2})$ is pairwise compact, by Proposition 2, we have $\mathcal{R}_{\Box}[\mathcal{W}]$ is pairwise compact.
Now we verify the condition $(ii)$ in the object part of Definition 10. Since $\{\langle a\rangle:a\in\mathcal{A}\}\in\beta_{1}$ and $\{\langle T_{1}(a)\rightarrow 0\rangle:a\in\mathcal{A}\}\in\beta_{2}$ are the basis for the topologies $\tau_{1}$ and $\tau_{2}$, respectively, so we show that for each $a\in\mathcal{A}$, $\langle\mathcal{R}_{\Box}\rangle(\langle a\rangle)\in\beta_{1}$ and $[\mathcal{R}_{\Box}]\langle a\rangle\in\beta_{1}$. We see that
@@
We show that $([\mathcal{R}_{\Box}]\langle T_{1}(a)\rightarrow 0\rangle)^{c}$ is $\tau_{1}$-open and $\tau_{2}$-closed. Let $\mathcal{U}\in([\mathcal{R}_{\Box}]\langle T_{1}(a)\rightarrow 0\rangle)^{c}$. Then $\mathcal{R}_{\Box}[\mathcal{U}]\not\subset\langle T_{1}(a)\rightarrow 0\rangle$. It is easy to see that $\exists$ $\tau_{1}$-open set $\langle\Box T_{1}(a)\rangle$ such that $\mathcal{U}\in\langle\Box T_{1}(a)\rangle$. Let $\mathcal{E}\in\langle\Box T_{1}(a)\rangle$. Then $\mathcal{E}(\Box T_{1}(a))=1$. Using the Kripke condition we have $1=\mathcal{E}(\Box T_{1}(a))=\bigwedge\{\mathcal{U}(T_{1}(a)):\mathcal{E}\mathcal{R}_{\Box}\mathcal{U}\}$. According to Lemma 2, $\mathcal{U}(T_{1}(a))$ is either 00 or $1$. Henceforth, for all $\mathcal{U}\in HOM_{\mathcal{VA}_{\mathcal{L}}}(\mathcal{A},\mathcal{L})$ with $\mathcal{E}\mathcal{R}_{\Box}\mathcal{U}$ we have $\mathcal{U}(T_{1}(a))=1$. As a result, $\mathcal{R}_{\Box}[\mathcal{E}]\not\subset\langle T_{1}(a)\rightarrow 0\rangle$ i.e., $\mathcal{E}\in([\mathcal{R}_{\Box}]\langle T_{1}(a)\rightarrow 0\rangle)^{c}$. Henceforth, $\mathcal{U}\in\langle\Box T_{1}(a)\rangle\subset([\mathcal{R}_{\Box}]\langle T_{1}(a)\rightarrow 0\rangle)^{c}$. Therefore, $[\mathcal{R}_{\Box}]\langle T_{1}(a)\rightarrow 0\rangle)^{c}$ is $\tau_{1}$-open i.e., $\langle\mathcal{R}_{\Box}\rangle\langle a\rangle$ is $\tau_{1}$-open.
@@
Next, we show that $[\mathcal{R}_{\Box}]\langle a\rangle\in\beta_{1}$. We see that

### v3

@@ Abstract
Fitting’s Heyting-valued logic and Heyting-valued modal logic have already been studied from an algebraic viewpoint. In addition to algebraic axiomatizations with the completeness of Fitting’s Heyting-valued logic and Heyting-valued modal logic, both topological and coalgebraic dualities have also been developed for algebras of Fitting’s Heyting-valued modal logic. Bitopological methods have recently been employed to investigate duality for Fitting’s Heyting-valued logic. However, the concepts of bitopology and bi-Vietoris coalgebras are conspicuously absent from the development of dualities for Fitting’s many-valued modal logic. With this study, we try to bridge that gap. The main results are bitopological and coalgebraic duality for Fitting’s many-valued modal logic. We develop a bitopological duality for algebras of Fitting’s Heyting-valued modal logic by extending known bitopological duality for Fitting’s non-modal logic. To develop coalgebraic duality, we adapt Lauridsen’s bi-Vietoris construction from the category of pairwise Stone spaces to the category $PBS_{\mathcal{L}}$ of $\mathcal{L}$-valued (with $\mathcal{L}$ a bounded finite distributive lattice, i.e., a Heyting algebra) pairwise Boolean spaces by incorporating a structure map, and from this obtain the $\mathcal{L}$-biVietoris functor. Finally, we establish dual equivalence between coalgebras for the $\mathcal{L}$-biVietoris functor and algebras of Fitting’s $\mathcal{L}$-valued modal logic. As a result, we conclude that Fitting’s Heyting-valued modal logic is sound and complete with respect to the coalgebras of $\mathcal{L}$-biVietoris functor. We also apply this coalgebraic approach to the bitopological duality to show the existence of cofree and final coalgebras, and to establish a Hennessy-Milner property.
@@ 1 Introduction
The algebraic axiomatization of a modified version of Fitting’s Heyting-valued modal logic has already been addressed in [12]. Maruyama [13] proposed Jónsson-Tarski topological duality (see [18, 19, 20]) for $\mathcal{L}$-$\mathcal{ML}$-algebras (algebras of Fitting’s Heyting-valued modal logic), while [14] developed both topological and coalgebraic dualities within a natural duality and coalgebra framework for general algebraic structures encompassing algebras of Fitting’s many-valued modal logic.
@@
Although these existing strands independently address substantial aspects of the field, Maruyama [14] offers a natural duality and coalgebraic framework sufficiently broad to incorporate algebras of Fitting’s many-valued modal logic, Priestley-based hyperspace semantics for positive modal logics have been integrated by Bezhanishvili et al. [8], and Lauridsen [7] presents a bi-Vietoris construction and coalgebraic completeness within the Pairwise Stone space context- a considerable gap persists. There is no existing literature that defines a duality of many-valued modal logic based on the techniques of bitopology. It is important to mention that there is a conspicuous lack of an explicit coalgebraic semantics based on a bitopological framework of many-valued modal logic. Furthermore, it does not provide a formal demonstration that this coalgebraic category is dual to the category of algebras of Fitting-style many-valued modal logic.
This paper fills that gap. We integrate these methodologies to develop bitopological and coalgebraic dualities for $\mathcal{L}$-$\mathcal{ML}$-algebras, where $\mathcal{L}$ is a semi-primal algebra having a bounded lattice reduct. Our general theory extends both the Jónsson-Tarski duality and the Abramsky-Kupke-Kurz-Venema coalgebraic duality [1, 21] in the setting of bitopological language.
We aim to develop a bitopological duality for algebras of Fitting’s Heyting-valued modal logic by setting up a notion of $PRBS_{\mathcal{L}}$ as a category of $\mathcal{L}$-valued pairwise Boolean spaces equipped with a structure map and a compatible binary relation. As a result, the natural duality theory and Jónsson-Tarski duality for modal algebras are expanded within the framework of bitopological languages, thereby generalizing the bitopological duality for Fitting’s $\mathcal{L}$-valued logic [15] to the modal case.
Building upon the bitopological duality for Fitting’s $\mathcal{L}$-valued modal logic, we derive a coalgebraic semantics for Fitting’s many-valued modal logic. The adaptation of Lauridsen’s [7] bi-Vietoris functor on the category of pairwise Stone spaces serves as the cornerstone of this development. We then aim to make this machinery fit the many-valued context by defining an endofunctor, denoted by $V_{\mathcal{L}}^{bi}$ on the category $PBS_{\mathcal{L}}$ of $\mathcal{L}$-valued pairwise Boolean spaces endowed with a structure map. This ensures the functor $V_{\mathcal{L}}^{bi}$ preserves the $\mathcal{L}$-valued structure that is necessary in handling the many-valued semantics. We do not claim that the bi-Vietoris mechanism is novel; rather, we adapt it in a many-valued environment and change the algebraic semantics to coalgebraic semantics in bitopological setting by establishing the coalgebraic duality theory.
This study provides the first coherent bitopological and coalgebraic structure of Fitting’s many-valued modal logic. Our main contributions are:
  • A bitopological duality between the category of algebras of Fitting’s many-valued modal logic and a category $PRBS_{\mathcal{L}}$ of $\mathcal{L}$ valued relational pairwise Boolean spaces.
  • A coalgebraic semantics based on a specified $\mathcal{L}$-biVietoris endofunctor on the category $PBS_{\mathcal{L}}$, revealing that the category of coalgebras of this $\mathcal{L}$-biVietoris endofunctor is isomorphic to $PRBS_{\mathcal{L}}$. This combination of algebra, bitopology, and coalgebra provides us with new, valuable insights for Fitting’s many-valued modal logic, including soundness, completeness, the Hennessy-Milner property for image-finite models, and the existence of final and cofree coalgebras.
Significance and connection to the existing dualities:
The framework of Maruyama’s study [14] provides an extensive natural duality and coalgebraic structure that encompasses the algebras of Fitting’s many-valued modal logic. Our findings agree with that paradigm but with a bitopological improvement that is specifically tailored to the multi-valued setting. In the case that we ignore the second topology and change the structure map in the standard topological setting, our framework is then the one discussed by Maruyama [14]. Consequently, the dualities proposed in our work constitute a conservative and structured enhancement of that general setting.
Maruyama’s work [13] establishes the Jónsson-Tarski duality for algebras of Fitting’s $\mathcal{L}$-valued modal logic within a standard topological framework. Our study adapts those methods to models characterised as $\mathcal{L}$-valued relational pairwise Boolean spaces, i.e., pairwise Boolean spaces equipped with a subalgebra-indexed meet-preserving structure map and a binary relation that is compatible with both the topologies and the structure map, thus developing a bitopological duality for the modal case. This methodology does not employ Vietoris machinery; rather, it is a direct adaptation of Jónsson-Tarski and natural duality to bitopological languages.
In [7], Lauridsen developed the bi-Vietoris construction and coalgebraic completeness for positive modal logic within the framework of pairwise Stone spaces. Our study (see Section 4) lifts that bi-Vietoris construction to the many-valued context: we define a $\mathcal{L}$-biVietoris functor on the category $PBS_{\mathcal{L}}$ of $\mathcal{L}$-valued pairwise Boolean spaces endowed with a structure map. The $\mathcal{L}$-biVietoris functor preserves the $\mathcal{L}$-valued structure, and its coalgebras coincide with our $\mathcal{L}$-valued relational model. Thus, our adaptation is non-trivial.
If we consider $\mathcal{L}=\mathbf{2}$, the two-element Boolean algebra, our framework reduces to the classical case: the structure map becomes trivial and on the canonical dual spaces, $\mathfrak{G}(\mathcal{A})=(HOM_{\mathcal{VA}_{\mathcal{L}}}(\mathcal{A},\mathcal{L}),\tau_{1},\tau_{2},\alpha_{\mathcal{A}})$ that are obtained by the duality functor (cf. Section 3), both the topologies $\tau_{1}$ and $\tau_{2}$ have the same basis, i.e., $\tau_{1}=\tau_{2}$. As a result, the category $PBS_{\mathbf{2}}$ matches the category of classical Stone spaces. Thus, in the two-valued case, our $\mathcal{L}$-biVietoris functor reflects the Lauridsen’s bi-Vietoris functor on pairwise Stone spaces with equal topologies.
Assumptions and Scope
Through out this study, we assume $\mathcal{L}$ as a truth-value set which is bounded finite distributive lattice, i.e., a Heyting algebra. We work with pairwise Boolean spaces equipped with a subalgebra-indexed meet preserving structure map and algebras of Fitting’s $\mathcal{L}$-valued modal logic.
Assuming this, we modify bi-Vietoris construction of Lauridsen to define the concept of $\mathcal{L}$-biVietoris functor $V_{\mathcal{L}}^{bi}$, work out both bitopological and coalgebraic dulity theorems for $\mathcal{L}$-valued modal logic, and generate the fundamental consequences: soundness and completeness, existence of final, cofree coalgebras, and Hennessy-Milner property on image-finite coalgebras of our adapted bi-Vietoris functor $V_{\mathcal{L}}^{bi}$.
@@
In Section 2, we review the fundamental notions of bitopological spaces and algebras of Fitting’s Heyting-valued logic, denoted as $\mathcal{L}$-$\mathcal{VL}$-algebras, and we also discuss the bitopological duality theorem for $\mathcal{L}$-$\mathcal{VL}$-algebras. Section 3 contains our first main result: a bitopological duality for Fitting’s Heyting-valued modal logic. Section 4 extends Lauridsen’s bi-Vietoris construction from the category of pairwise Stone spaces to the category of $\mathcal{L}$-valued pairwise Boolean spaces, $PBS_{\mathcal{L}}$, resulting in an endofunctor $V_{\mathcal{L}}^{bi}$ on the category $PBS_{\mathcal{L}}$ that preserves the many-valued structure. We identify the category of $\mathcal{L}$-valued relational pairwise Boolean spaces $PRBS_{\mathcal{L}}$, with the category $COALG(V_{\mathcal{L}}^{bi})$ of coalgebras of the functor $V_{\mathcal{L}}^{bi}$, and by combining this with the result in Section 3, we obtain the coalgebraic duality for Fitting’s Heyting-valued modal logic and this yields soundness and completeness of Fitting’s many-valued modal logic with respect to $V_{\mathcal{L}}^{bi}$-coalgebras. In Section 5, we provide the applications of our coalgebraic approach to the bitopological duality. The fundamental applications are:
  • A demonstration of the Hennessy-Milnar property for $V_{\mathcal{L}}^{bi}$-coalgebras.
  • The existence of cofree coalgebras, and final coalgebras.
In Section 6, we conclude the paper by indicating the limitations of our study and provide some possible avenues of future research.
@@ 2.2 Fitting’s Heyting-valued logics
Fitting [11] proposed $\mathcal{L}$-valued logics and $\mathcal{L}$-valued modal logics for a finite distributive lattice $\mathcal{L}$(i.e., $\mathcal{L}$ is a Heyting algebra) in 1991. Maruyama [12] introduced algebraic axiomatization of Fitting’s logics. In [12] the author studied Fitting’s Heyting-valued logic and Heyting-valued modal logic without regard for fuzzy truth constants other than 0 and $1$, and added a new operation $T_{\ell}(-)$, $\ell\in\mathcal{L}$. From a logical perspective, $T_{\ell}(p)$ infers that the truth value of a proposition $p$ is $\ell$. The operations of $\mathcal{L}$-valued logic, denoted by $\mathcal{L}$-$\mathcal{VL}$, are $\vee,\wedge,\rightarrow,0,1$ and $T_{\ell}(-)$, $\ell\in\mathcal{L}$, where $\vee,\wedge,\rightarrow$ are binary operations, 0 and $1$ are nullary operations and $T_{\ell}$ is a unary operation. For $\ell_{1},\ell_{2}\in\mathcal{L}$, $\ell_{1}\rightarrow\ell_{2}$ means the pseudo-complement of $\ell_{1}$ relative to $\ell_{2}$.
@@
[…]
@@
[…]
@@
The class of all $\mathcal{L}$-$\mathcal{VL}$-algebras forms a variety (in the sense of universal algebra). If $\mathcal{L}=\{0,1\}$, then $\mathcal{L}$-$\mathcal{VL}$-algebras becomes Boolean algebras.
@@ 2.3 Bitopological duality for Fitting’s Heyting-valued logic
    For an object $(\mathcal{B},\alpha_{\mathcal{B}})$ in $PBS_{\mathcal{L}}$, define $\mathfrak{F}(\mathcal{B},\alpha_{\mathcal{B}})=(HOM_{PBS_{\mathcal{L}}}((\mathcal{B},\alpha_{\mathcal{B}}),(\mathcal{L},\alpha_{\mathcal{L}})),\vee,\wedge,\rightarrow,T_{\ell}(\ell\in\mathcal{L}),0,1)$, where $\vee,\wedge,\rightarrow,T_{\ell}(\ell\in\mathcal{L}),0,1$ are pointwise operations on the set $HOM_{PBS_{\mathcal{L}}}((\mathcal{B},\alpha_{\mathcal{B}}),(\mathcal{L},\alpha_{\mathcal{L}}))$. The operations 0 and $1$ are regarded as constant functions, with 0 and $1$ being their respective values.
@@ 3 Bitopological duality for Fitting’s many-valued modal logic
This section extends modal natural duality framework developed by Maruyama in [14] to the bitopological language. On the semantic side we work with the category $PRBS_{\mathcal{L}}$ of $\mathcal{L}$-valued relational pairwise Boolean spaces. We establish $\mathcal{MA}_{\mathcal{L}}\simeq PRBS_{\mathcal{L}}$. When $\mathcal{L}=\mathbf{2}$, the two-element Boolean algebra, our duality shows the Jónsson-Tarski duality for modal algebras.
@@
$\mathcal{R}]C=\{p\in P:\mathcal{R}[p]\subseteq C\}$ and $\langle\mathcal{R}\rangle C=\{p\in P:\mathcal{R}[p]\cap C\neq\emptyset\}$.
@@ 3.1 Category
      $\forall\mathcal{C}\in\beta_{1}$, $\mathcal{R}]\mathcal{C},\langle\mathcal{R}\rangle\mathcal{C}\in\beta_{1}$;
@@
We see that $\mathcal{R}]Q^{c}=(\langle\mathcal{R}\rangle Q)^{c}$, and $\langle\mathcal{R}\rangle Q^{c}=([\mathcal{R}]Q)^{c}$. Since $\beta_{2}=\{Q^{c}:Q\in\beta_{1}\}$, hence if the relation $\mathcal{R}$ satisfies condition $(ii)$ that is given in the object part of Definition 10, then $\mathcal{R}]Q,\langle\mathcal{R}\rangle Q\in\beta_{2}$, $\forall Q\in\beta_{2}$.
@@ 3.2 Functors
Definition 9 shows that $(HOM_{\mathcal{VA}_{\mathcal{L}}}(\mathcal{A},\mathcal{L}),\tau_{1},\tau_{2},\alpha_{\mathcal{A}})$ is an object in $PBS_{\mathcal{L}}$. Thus, it suffices to demonstrate $\mathcal{R}_{\Box}$ satisfies the requirements given in the object section of Definition 10. First, we ensure that for any $\mathcal{W}\in HOM_{\mathcal{VA}_{\mathcal{L}}}(\mathcal{A},\mathcal{L})$, $\mathcal{R}_{\Box}[\mathcal{W}]\in\delta_{1}\cup\delta_{2}$. Let $\mathcal{U}\not\in\mathcal{R}_{\Box}[\mathcal{W}]$. Then by Definition 6, there is an element $a\in\mathcal{A}$ such that there is $L_{1}\in\mathcal{L}$, for which $\mathcal{W}(\Box a)\geq L_{1}$ but $\mathcal{U}(a)\not\geq L_{1}$. It follows that $\mathcal{U}\in\langle\neg U_{L_{1}}(a)\rangle\in\tau_{2}$ and $\mathcal{R}_{\Box}[\mathcal{W}]\cap\langle\neg U_{L_{1}}(a)\rangle=\emptyset$ i.e., $\langle\neg U_{L_{1}}(a)\rangle\subseteq(\mathcal{R}_{\Box}[\mathcal{W}])^{c}$. Hence, $\mathcal{U}\not\in\overline{\mathcal{R}_{\Box}[\mathcal{W}]}^{\tau_{2}}$, where $\overline{\mathcal{R}_{\Box}[\mathcal{W}]}^{\tau_{2}}$ denotes the closure of $\mathcal{R}_{\Box}[\mathcal{W}]$ in $(HOM_{\mathcal{VA}_{\mathcal{L}}}(\mathcal{A},\mathcal{L}),\tau_{2})$ . Equivalently, we have $\overline{\mathcal{R}_{\Box}[\mathcal{W}]}^{\tau_{2}}\subset\mathcal{R}_{\Box}[\mathcal{W}]$. Therefore, $\mathcal{R}_{\Box}[\mathcal{W}]$ is $\tau_{2}$-closed. Since $(HOM_{\mathcal{VA}_{\mathcal{L}}}(\mathcal{A},\mathcal{L}),\tau_{1},\tau_{2})$ is pairwise compact, by Proposition 2, we have $\mathcal{R}_{\Box}[\mathcal{W}]$ is pairwise compact.
Now we verify the condition $(ii)$ in the object part of Definition 10. Since $\{\langle a\rangle:a\in\mathcal{A}\}\in\beta_{1}$ and $\{\langle T_{1}(a)\rightarrow 0\rangle:a\in\mathcal{A}\}\in\beta_{2}$ are the basis for the topologies $\tau_{1}$ and $\tau_{2}$, respectively, so we show that for each $a\in\mathcal{A}$, $\langle\mathcal{R}_{\Box}\rangle(\langle a\rangle)\in\beta_{1}$ and $\mathcal{R}_{\Box}]\langle a\rangle\in\beta_{1}$. We see that
@@
We show that $([\mathcal{R}_{\Box}]\langle T_{1}(a)\rightarrow 0\rangle)^{c}$ is $\tau_{1}$-open and $\tau_{2}$-closed. Let $\mathcal{U}\in([\mathcal{R}_{\Box}]\langle T_{1}(a)\rightarrow 0\rangle)^{c}$. Then $\mathcal{R}_{\Box}[\mathcal{U}]\not\subset\langle T_{1}(a)\rightarrow 0\rangle$. It is easy to see that $\exists$ $\tau_{1}$-open set $\langle\Box T_{1}(a)\rangle$ such that $\mathcal{U}\in\langle\Box T_{1}(a)\rangle$. Let $\mathcal{E}\in\langle\Box T_{1}(a)\rangle$. Then $\mathcal{E}(\Box T_{1}(a))=1$. Using the Kripke condition we have $1=\mathcal{E}(\Box T_{1}(a))=\bigwedge\{\mathcal{U}(T_{1}(a)):\mathcal{E}\mathcal{R}_{\Box}\mathcal{U}\}$. According to Lemma 2, $\mathcal{U}(T_{1}(a))$ is either 0 or $1$. Henceforth, for all $\mathcal{U}\in HOM_{\mathcal{VA}_{\mathcal{L}}}(\mathcal{A},\mathcal{L})$ with $\mathcal{E}\mathcal{R}_{\Box}\mathcal{U}$ we have $\mathcal{U}(T_{1}(a))=1$. As a result, $\mathcal{R}_{\Box}[\mathcal{E}]\not\subset\langle T_{1}(a)\rightarrow 0\rangle$ i.e., $\mathcal{E}\in([\mathcal{R}_{\Box}]\langle T_{1}(a)\rightarrow 0\rangle)^{c}$. Henceforth, $\mathcal{U}\in\langle\Box T_{1}(a)\rangle\subset([\mathcal{R}_{\Box}]\langle T_{1}(a)\rightarrow 0\rangle)^{c}$. Therefore, $\mathcal{R}_{\Box}]\langle T_{1}(a)\rightarrow 0\rangle)^{c}$ is $\tau_{1}$-open i.e., $\langle\mathcal{R}_{\Box}\rangle\langle a\rangle$ is $\tau_{1}$-open.
@@
Next, we show that $\mathcal{R}_{\Box}]\langle a\rangle\in\beta_{1}$. We see that